OpenStudy (anonymous):

Solve \[(x/x-2)+(3/x+3)=(-14/x^2+x-6)\]

OpenStudy (lgbasallote):

if i tell you that x^2 + x - 6 = (x-2)(x+3) will that help?

OpenStudy (anonymous):

Parentheses in the right places? Or does it look like this: \[\frac{x}{x-2}+\frac{3}{x+3}=\frac{-14}{x^2+x-6}\]

OpenStudy (anonymous):

it looks that that

OpenStudy (mertsj):

Multiply each term by the common denominator which is (x-2)(x+3)

OpenStudy (mertsj):

x(x+3)+3(x-2)=-14

OpenStudy (mertsj):

Take it from there.

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

so x-2 i only multiply the x+3

OpenStudy (anonymous):

i only multiply the missing ones?

OpenStudy (anonymous):

When you multiply every term by (x-2)(x+3), what do you get?

OpenStudy (anonymous):

But aren't you supposed to cancel out terms that are already there?

OpenStudy (anonymous):

Right :) . What do you get after you do that?

OpenStudy (mertsj):

\[\frac{x}{x-2}\times(x-2)(x+3)+\frac{3}{x+3}\times(x-2)(x+3)=\frac{-14}{(x-2)(x+3})\times(x-2)(x+3)\]

OpenStudy (mertsj):

\[x(x+3)+3(x-2)=-14\]